# supplementary_materials

### v1

@@ Appendix C Estimation of $V_{DW}$
$U_{sc}(\mathbf{q})\rightarrow U_{0}=\frac{e^{2}d_{g}}{2\epsilon_{0}\epsilon}$ (14)
@@
We expect this approximation to be reasonable at small densities, therefore small Fermi momenta, such that the typical momenta are much smaller than $1/d_{g}$. Using the Hubbard interaction $U_{0}$, we can now compute the bubble and ladder sums (following Ref. Berk and Schrieffer (1966); Scalapino et al. (1986)) to obtain the “paramagnon” mediated interaction:
$\frac{1}{A}\sum_{\mathbf{k},\mathbf{k}^{\prime},\mathbf{q}}V_{DW}(\mathbf{k},\mathbf{k}^{\prime},\mathbf{q})f^{\dagger}_{v,\mathbf{k}+\mathbf{q}}f_{c,\mathbf{k}}f^{\dagger}_{c,\mathbf{k}^{\prime}}f_{v,\mathbf{k}^{\prime}+\mathbf{q}}$ (15)
$V_{DW}(\mathbf{k},\mathbf{k}^{\prime},\mathbf{q})$ $=-\frac{U_{0}^{2}\Pi_{cv}(\mathbf{q})}{1+U_{0}\Pi_{cv}(\mathbf{q})}+\frac{U_{0}}{1-U_{0}(\Pi_{cc}(\mathbf{k}-\mathbf{k}^{\prime})+\Pi_{vv}(\mathbf{k}-\mathbf{k}^{\prime}))}$ (16)
@@
where $\Pi_{cv}$ is the inter-band susceptibility and $\Pi_{cc/vv}$ are the corresponding intra-band susceptibilities. The first term is from the ladder sum and the second is due to the bubble sum (we have included the bare Hubbard interaction in the second term). The first term dominates near the DW-QCP and gets peaked near $\mathbf{q}=\pm\mathbf{Q}$, therefore, we ignore the second term in our analysis. We fit this interaction to the form specified in Eq. 8 of the main text to extract fit parameters $g,m_{b}$ and $\kappa$, as a function of $n$ and $V_{b}$. The fit parameters used in our calculations are plotted in Fig. 7.
The density wave instability onsets when $U_{0}\Pi_{cv}(\mathbf{q})=-1$ at momentum $\mathbf{q}$. We present in Fig. 8 the behavior of $U_{0}\Pi_{cv}(\mathbf{q})$ in the symmetric phase as a function of $\mathbf{q}$ along the $x$ and $y$ directions. As seen, the leading instability is for $\mathbf{q}\sim\hat{x}$.
@@
Figure 7: Top left: The criterion for the X-DW phase transition within the Hubbard interaction approximation. Yellow and blue are respectively the symmetric ($\Pi_{cv}(\mathbf{Q})U_{0}>-1$) and the X-DW ($\Pi_{cv}(\mathbf{Q})U_{0}<-1$) phases. The other panels plot the interaction parameters $g,\kappa,m_{b}$ defined in $V_{DW}$ as a function of phase space parameters. Parameters used are: $\epsilon=18$, $199\times 199$ momentum mesh.
@@
Figure 8: $U_{0}\Pi_{cv}$ as a function of momentum. As can be seen, $U_{0}\Pi_{cv}$ along $\hat{x}$-direction is able to satisfy the density wave instability condition at smaller interaction strengths.
Appendix D Goldstone Mode Mediated Interactions
@@
The two X-DW phases discussed in this work have gapless Goldstone modes due to breaking of continuous translation symmetry and the $U(1)$ interlayer charge conservation symmetry. The goal of this section is to study whether the interactions mediated by them can also lead to superconductivity. Our analysis will be within the Hubbard approximation.
D.1 Collinear X-DW
@@
[…]
Here $\chi_{\mathbf{r}}=\chi_{0}\cos\mathbf{Q}\cdot\mathbf{r}+\delta\chi_{\mathbf{r}}=\chi_{0}\cos{\mathbf{Q}\cdot\mathbf{r}}+\delta\chi_{-}(\mathbf{r})e^{-i\mathbf{Q}\cdot\mathbf{r}}+\delta\chi_{+}(\mathbf{r})e^{i\mathbf{Q}\cdot\mathbf{r}}$ where the collinear X-DW has a periodicity given by Q, $\chi_{0}$ is the uniform mean field exciton condensate, and $\delta\chi_{\mathbf{r}}$ is the fluctuation. For Hubbard interactions $U_{0}$, the corresponding momentum space mean-field order parameter is $\mathbf{k}$-independent, and is self-consistently determined by solving the mean-field Hamiltonian. We now focus on the fluctuating part of Eq. 17 and rewrite it in momentum space as,
[…]
@@
$H_{fluc.}=\frac{A}{2}\sum_{\mathbf{q}}\delta\chi_{i,y}(-\mathbf{q})\left(\frac{2}{U_{0}}\delta_{ij}-\Pi_{ij}(\mathbf{q},\Omega=0)\right)\delta\chi_{j,y}(\mathbf{q}).$ (19)
Here $i,j$ run over the two fluctuations $\pm$, $A$ is the sample area, and $\delta_{ij}$ is the Kronecker delta function. $\Pi_{ij}$ is calculated using the mean-field quasiparticle states. It is evident from the first term in Eq. 19 that we are employing the Hubbard approximation. The polarization bubble is given by,
$\Pi_{ij}(\mathbf{q},\Omega)=-\int\frac{d^{2}kd\omega}{(2\pi)^{3}}\text{Tr}\left(\sigma_{i,y}G(\mathbf{k}+\mathbf{q},\omega+\Omega)\sigma_{j,y}G(\mathbf{k},\omega)\right),$ (20)
@@
$H_{\text{fluc}}=\frac{1}{4}\sum_{\mathbf{q}}\begin{pmatrix}\delta\chi_{-,y}(-\mathbf{q})&\delta\chi_{+,y}(-\mathbf{q})\end{pmatrix}\left((\rho^{x}_{s}+\rho^{x}_{p})q_{x}^{2}+(\rho^{y}_{s}+\rho^{y}_{p})q_{y}^{2}+\sigma_{x}((\rho^{x}_{s}-\rho^{x}_{p})q_{x}^{2}+(\rho^{y}_{s}-\rho^{y}_{p})q_{y}^{2})\right)\begin{pmatrix}\delta\chi_{-,y}(\mathbf{q})\\ \delta\chi_{+,y}(\mathbf{q})\end{pmatrix}$ (21)
@@
Let us now calculate the coupling of these modes to the fermionic quasiparticles in the collinear X-DW phase. We define a Brillouin zone (BZ) along $\hat{x}$ direction due to the collinear X-DW ordering: $k_{x}\in[-Q,Q)$. For generic $(k_{F,c}$, $k_{F,v})$, the Fermi surface can be complicated due to bands originating in multiple zones that cross the Fermi level. Therefore, we focus on parameter points $(n,V_{b})$ such that $k_{F,c,x}<Q$, which simplifies the Fermi surface to arise from only three bands: $f_{c,\mathbf{k}}$, $f_{v,\mathbf{k}+\mathbf{Q}}$ and $f_{c,\mathbf{k}-\mathbf{Q}}$, where $\mathbf{k}$ is defined in the BZ. For Hubbard interactions, the mean-field Hamiltonian is
$H_{mf}=\sum_{\mathbf{k}}\begin{pmatrix}f^{\dagger}_{c,\mathbf{k}}&f^{\dagger}_{v,\mathbf{k}+\mathbf{Q}}&f^{\dagger}_{v,\mathbf{k}-\mathbf{Q}}\end{pmatrix}\begin{pmatrix}\xi_{c,\mathbf{k}}&-\chi_{+}&-\chi_{-}\\ -\chi^{*}_{+}&\xi_{v,\mathbf{k}+\mathbf{Q}}&0\\ -\chi^{*}_{-}&0&\xi_{v,\mathbf{k}-\mathbf{Q}}\end{pmatrix}\begin{pmatrix}f_{c,\mathbf{k}}\\ f_{v,\mathbf{k}+\mathbf{Q}}\\ f_{v,\mathbf{k}-\mathbf{Q}}\end{pmatrix}$ (22)
@@
We numerically obtain real self-consistent solutions such that $\chi_{+}=\chi_{-}$, and find the wavefunction $\psi_{\mathbf{k}}$ (defined in the 3-band basis) of the low-energy band $f_{\mathbf{k}}$ that forms the Fermi surface with dispersion $\epsilon_{\mathbf{k}}$. We can then obtain the fermion-Goldstone mode coupling vertex $\alpha_{G}$:
$\alpha_{s}(\mathbf{k}^{\prime},\mathbf{k})$ $=\psi_{\mathbf{k}^{\prime}}^{*}\begin{pmatrix}0&-1&-1\\ 1&0&0\\ 1&0&0\end{pmatrix}\psi_{\mathbf{k}}$ (23)
$\alpha_{p}(\mathbf{k}^{\prime},\mathbf{k})$ $=\psi_{\mathbf{k}^{\prime}}^{*}\begin{pmatrix}0&-1&1\\ 1&0&0\\ -1&0&0\end{pmatrix}\psi_{\mathbf{k}}$ (24)
The interaction term that is mediated by the Goldstone modes is:
$\frac{1}{2A\rho_{G}}\sum_{\mathbf{k},\mathbf{k}^{\prime},\mathbf{q}}f^{\dagger}_{\mathbf{k}+\mathbf{q}}f_{\mathbf{k}}f^{\dagger}_{\mathbf{k}^{\prime}-\mathbf{q}}f_{\mathbf{k}^{\prime}}\frac{\alpha_{G}(\mathbf{k}+\mathbf{q},\mathbf{k})\alpha_{G}(\mathbf{k}^{\prime}-\mathbf{q},\mathbf{k}^{\prime})}{q_{x}^{2}+\kappa q_{y}^{2}}.$ (25)
@@
We note that the strength of the interaction approaches a finite limit as $\mathbf{q}\rightarrow 0$, analogous to the case of typical Goldstone modes such as acoustic phonons and gapless spin waves Watanabe and Vishwanath (2014), since $\alpha(\mathbf{k},\mathbf{q})$ is linear in momentum for small momentum. Let us define the Goldstone mediated interactions strength:
$V_{G}(\mathbf{k},\mathbf{k}^{\prime},\mathbf{q})=\frac{1}{\rho_{G}}\frac{\alpha_{G}(\mathbf{k},\mathbf{q})\alpha_{G}(\mathbf{k}^{\prime},-\mathbf{q})}{q_{x}^{2}+\kappa_{G}q_{y}^{2}}.$ (26)
@@
$\frac{1}{2A}\sum_{\mathbf{k},\mathbf{l}}f^{\dagger}_{\mathbf{l}}f^{\dagger}_{-\mathbf{l}}f_{-\mathbf{k}}f_{\mathbf{k}}V_{G}(\mathbf{k},-\mathbf{k},\mathbf{l}-\mathbf{k}).$ (27)
Let us define $V_{BCS}(\mathbf{k},\mathbf{l})=V_{G}(\mathbf{k},-\mathbf{k},\mathbf{l}-\mathbf{k})/A$. We plot $V_{BCS}$ this interaction for $l_{x}\rightarrow k_{x},l_{y}=k_{y}$ in Fig. 9, to demonstrate the attractive nature of the interactions. This motivates us to ask whether the Goldstone mode mediated interaction can lead to superconductivity. Decoupling the interaction term in the BCS channel, we get the following BdG Hamiltonian.
$H_{\text{BdG}}=\sum_{\mathbf{k}}\begin{pmatrix}f^{\dagger}_{\mathbf{k}}&f_{-\mathbf{k}}\end{pmatrix}\begin{pmatrix}\epsilon_{\mathbf{k}}&\Delta_{BCS}(\mathbf{k})\\ \Delta^{*}_{BCS}(\mathbf{k})&-\epsilon_{-\mathbf{k}}\end{pmatrix}\begin{pmatrix}f_{\mathbf{k}}\\ f^{\dagger}_{-\mathbf{k}}\end{pmatrix}$ (28)
@@
$\Delta_{BCS}(\mathbf{k})=\sum_{\mathbf{l}}V_{BCS}(\mathbf{k},\mathbf{l})\langle f_{-,-\mathbf{l}}f_{-,\mathbf{l}}\rangle$ (29)
$\epsilon_{\mathbf{k}}=\epsilon_{-\mathbf{k}}$ due to the $C_{2z}$ symmetry. For a range of parameters, we find self-consistent superconducting solutions, as plotted in Fig. 4. The gap is maximized in the region where the interactions are strongly attractive.
Figure 9: Left/Right panel: Phonon/superfluid mode mediated interaction matrix elements for ${\mathbf{l}}\rightarrow{\mathbf{k}}$ in Eq. 27. The parameters used are $n=-0.7\times 10^{12}$ cm-2, $V_{b}=12$ meV and a $59\times 59$ momentum mesh.
D.2 Spiral X-DW
For completness, we also examine the situation of Goldstone mediated superconductivity in the spiral X-DW. The Hubbard-Stratonovich field associated with the sprial X-DW is $\chi_{\mathbf{r}}=e^{i\textbf{Q}\cdot\mathbf{r}}(\chi+\delta\chi_{\mathbf{r}})$ where the spiral X-DW has a periodicity given by Q, $\chi$ is the uniform mean field exciton condensate, and $\delta\chi_{\mathbf{r}}$ is the fluctuation. We note that we have extracted out the fast oscillatory behavior (arising from Q) explicitly in the order parameter. The corresponding momentum space mean-field order parameter is self-consistently determined from the gap equation (which follows from Eq. 13),
$\chi=\frac{U_{0}}{A}\sum_{\mathbf{k}^{\prime}}\left(n_{F}(\epsilon_{-,\mathbf{k}})-n_{F}(\epsilon_{+,\mathbf{k}})\right)\frac{\chi}{2E_{\mathbf{k}^{\prime}}}.$ (30)

### v2

@@ III Estimation of $V_{DW}$
$U_{sc}(\mathbf{q})\rightarrow U_{0}=\frac{e^{2}d_{g}}{2\epsilon_{0}\epsilon}$ (A.12)
@@
We expect this approximation to be reasonable at small densities, therefore small Fermi momenta, such that the typical momenta are much smaller than $1/d_{g}$. Using the Hubbard interaction $U_{0}$, we can now compute the bubble and ladder sums (following Ref. [6, 55]) to obtain the “paramagnon” mediated interaction:
$\frac{1}{A}\sum_{\mathbf{k},\mathbf{k}^{\prime},\mathbf{q}}V_{DW}(\mathbf{k},\mathbf{k}^{\prime},\mathbf{q})f^{\dagger}_{v,\mathbf{k}+\mathbf{q}}f_{c,\mathbf{k}}f^{\dagger}_{c,\mathbf{k}^{\prime}}f_{v,\mathbf{k}^{\prime}+\mathbf{q}}$ (A.13)
$V_{DW}(\mathbf{k},\mathbf{k}^{\prime},\mathbf{q})$ $=-\frac{U_{0}^{2}\Pi_{cv}(\mathbf{q})}{1+U_{0}\Pi_{cv}(\mathbf{q})}+\frac{U_{0}}{1-U_{0}(\Pi_{cc}(\mathbf{k}-\mathbf{k}^{\prime})+\Pi_{vv}(\mathbf{k}-\mathbf{k}^{\prime}))}$ (A.14)
@@
where $\Pi_{cv}$ is the inter-band susceptibility and $\Pi_{cc/vv}$ are the corresponding intra-band susceptibilities. The first term is from the ladder sum and the second is due to the bubble sum (we have included the bare Hubbard interaction in the second term). The first term dominates near the DW-QCP and gets peaked near $\mathbf{q}=\pm\mathbf{Q}$, therefore, we ignore the second term in our analysis. We fit this interaction to the form specified in Eq. 8 of the main text to extract fit parameters $g,m_{b}$ and $\kappa$, as a function of $n$ and $V_{b}$. The fit parameters used in our calculations are plotted in Fig. A.2.
The density wave instability onsets when $U_{0}\Pi_{cv}(\mathbf{q})=-1$ at momentum $\mathbf{q}$. We present in Fig. A.3 the behavior of $U_{0}\Pi_{cv}(\mathbf{q})$ in the symmetric phase as a function of $\mathbf{q}$ along the $x$ and $y$ directions. As seen, the leading instability is for $\mathbf{q}\sim\hat{x}$. The $\hat{x}$ direction is picked because of the larger effective mass of the valence band along that direction. If, instead, the system exhibited continuous rotational symmetry, all directions will be equally susceptible. Therefore, at the phase transition, there will be a circle in momentum space on which the exciton fluctuations become gapless. We anticipate the enhanced fluctuations from these gapless modes will turn the transition first order. We will hence focus on the discrete $C_{2z}$ symmetric case where the exciton modes become gapless only at two points.
@@
Figure A.2: Top left: The criterion for the X-DW phase transition within the Hubbard interaction approximation. Yellow and blue are respectively the symmetric ($\Pi_{cv}(\mathbf{Q})U_{0}>-1$) and the X-DW ($\Pi_{cv}(\mathbf{Q})U_{0}<-1$) phases. The other panels plot the interaction parameters $g,\kappa,m_{b}$ defined in $V_{DW}$ as a function of phase space parameters. Parameters used are: $\epsilon=18$, $199\times 199$ momentum mesh.
@@
Figure A.3: $U_{0}\Pi_{cv}$ as a function of momentum. As can be seen, $U_{0}\Pi_{cv}$ along $\hat{x}-$direction is able to satisfy the density wave instability condition at smaller interaction strengths.
IV Goldstone Mode Mediated Interactions
@@
The two X-DW phases discussed in this work have gapless Goldstone modes due to breaking of continuous translation symmetry and the $U(1)$ interlayer charge conservation symmetry. The goal of this section is to study whether the interactions mediated by them can also lead to superconductivity. For analytical tractability of our analysis, we assume that the gate-screened Coulomb interactions become over-screened in the limit of small gate distances $d_{g}\ll 1/q$, leading to a momentum-independent Hubbard interaction.
IV.1 Collinear X-DW
@@
$H_{\chi-f}=-\sum_{\mathbf{r}}\chi_{\mathbf{r}}f_{v,\mathbf{r}}^{{\dagger}}f_{c,\mathbf{r}}+\text{h.c.}$ (A.15)
Here $\chi_{\mathbf{r}}=\chi_{0}\cos\mathbf{Q}\cdot\mathbf{r}+\delta\chi_{\mathbf{r}}=\chi_{0}\cos{\mathbf{Q}\cdot\mathbf{r}}+\delta\chi_{-}(\mathbf{r})e^{-i\mathbf{Q}\cdot\mathbf{r}}+\delta\chi_{+}(\mathbf{r})e^{i\mathbf{Q}\cdot\mathbf{r}}$ where the collinear X-DW has a periodicity given by Q, $\chi_{0}$ is the uniform mean field exciton condensate, and $\delta\chi_{\mathbf{r}}$ is the fluctuation. For Hubbard interactions $U_{0}$, the corresponding momentum space mean-field order parameter is $\mathbf{k}$-independent, and is self-consistently determined by solving the mean-field Hamiltonian. We now focus on the fluctuating part of Eq. A.15 and rewrite it in momentum space as,
$H_{\delta\chi-f}$ $=-\sum_{\mathbf{q},\mathbf{k}}\delta\chi_{-}({-\mathbf{q}})f_{c,\mathbf{k}}^{{\dagger}}f_{v,\mathbf{k}-\textbf{Q}+\mathbf{q}}+\delta\chi_{+}(-{\mathbf{q}})f_{c,\mathbf{k}}^{{\dagger}}f_{v,\mathbf{k}+\textbf{Q}+\mathbf{q}}+\text{h.c.}$ (A.16)
@@
$H_{fluc.}=\frac{A}{2}\sum_{\mathbf{q}}\delta\chi_{i,y}(-\mathbf{q})\left(\frac{2}{U_{0}}\delta_{ij}-\Pi_{ij}(\mathbf{q},\Omega=0)\right)\delta\chi_{j,y}(\mathbf{q}).$ (A.17)
Here $i,j$ run over the two fluctuations $\pm$, $A$ is the sample area, and $\delta_{ij}$ is the Kronecker delta function. $\Pi_{ij}$ is calculated using the mean-field quasiparticle states. It is evident from the first term in Eq. A.17 that we are employing the Hubbard approximation. The polarization bubble is given by,
$\Pi_{ij}(\mathbf{q},\Omega)=-\int\frac{d^{2}kd\omega}{(2\pi)^{3}}\text{Tr}\left(\sigma_{i,y}G(\mathbf{k}+\mathbf{q},\omega+\Omega)\sigma_{j,y}G(\mathbf{k},\omega)\right),$ (A.18)
@@
$H_{\text{fluc}}=\frac{1}{4}\sum_{\mathbf{q}}\begin{pmatrix}\delta\chi_{-,y}(-\mathbf{q})&\delta\chi_{+,y}(-\mathbf{q})\end{pmatrix}\left((\rho^{x}_{s}+\rho^{x}_{p})q_{x}^{2}+(\rho^{y}_{s}+\rho^{y}_{p})q_{y}^{2}+\sigma_{x}((\rho^{x}_{s}-\rho^{x}_{p})q_{x}^{2}+(\rho^{y}_{s}-\rho^{y}_{p})q_{y}^{2})\right)\begin{pmatrix}\delta\chi_{-,y}(\mathbf{q})\\ \delta\chi_{+,y}(\mathbf{q})\end{pmatrix}$ (A.19)
@@
Let us now calculate the coupling of these modes to the fermionic quasiparticles in the collinear X-DW phase. We define a Brillouin zone (BZ) along $\hat{x}$ direction due to the collinear X-DW ordering: $k_{x}\in[-Q,Q)$. For generic $(k_{F,c}$, $k_{F,v})$, the Fermi surface can be complicated due to bands originating in multiple zones that cross the Fermi level. Therefore, we focus on parameter points $(n,V_{b})$ such that $k_{F,c,x}<Q$, which simplifies the Fermi surface to arise from only three bands: $f_{c,\mathbf{k}}$, $f_{v,\mathbf{k}+\mathbf{Q}}$ and $f_{v,\mathbf{k}-\mathbf{Q}}$, where $\mathbf{k}$ is defined in the BZ. For Hubbard interactions, the mean-field Hamiltonian is
$H_{mf}=\sum_{\mathbf{k}}\begin{pmatrix}f^{\dagger}_{c,\mathbf{k}}&f^{\dagger}_{v,\mathbf{k}+\mathbf{Q}}&f^{\dagger}_{v,\mathbf{k}-\mathbf{Q}}\end{pmatrix}\begin{pmatrix}\xi_{c,\mathbf{k}}&-\chi_{+}&-\chi_{-}\\ -\chi^{*}_{+}&\xi_{v,\mathbf{k}+\mathbf{Q}}&0\\ -\chi^{*}_{-}&0&\xi_{v,\mathbf{k}-\mathbf{Q}}\end{pmatrix}\begin{pmatrix}f_{c,\mathbf{k}}\\ f_{v,\mathbf{k}+\mathbf{Q}}\\ f_{v,\mathbf{k}-\mathbf{Q}}\end{pmatrix}$ (A.20)
@@
We numerically obtain real self-consistent solutions such that $\chi_{+}=\chi_{-}$, and find the wavefunction $\psi_{\mathbf{k}}$ (defined in the 3-band basis) of the low-energy band $f_{\mathbf{k}}$ that forms the Fermi surface with dispersion $\epsilon_{\mathbf{k}}$. We can then obtain the fermion-Goldstone mode coupling between the low-energy fermions and the Goldstone modes:
$\alpha_{s}(\mathbf{k}^{\prime},\mathbf{k})$ $=\psi_{\mathbf{k}^{\prime}}^{*}\begin{pmatrix}0&-1&-1\\ 1&0&0\\ 1&0&0\end{pmatrix}\psi_{\mathbf{k}}$
$\alpha_{p}(\mathbf{k}^{\prime},\mathbf{k})$ $=\psi_{\mathbf{k}^{\prime}}^{*}\begin{pmatrix}0&-1&1\\ 1&0&0\\ -1&0&0\end{pmatrix}\psi_{\mathbf{k}}$
$H_{f-G}$ $=-i\int_{\mathbf{q},\mathbf{k}}\alpha_{G}(\mathbf{k}+\mathbf{q},\mathbf{k})\delta\chi_{G}(-\mathbf{q})f^{\dagger}_{\mathbf{k}+\mathbf{q}}f_{\mathbf{k}}.$ (A.21)
Writing $\psi_{\mathbf{k}}=\begin{bmatrix}a_{\mathbf{k}}&b_{\mathbf{k}}&c_{\mathbf{k}}\end{bmatrix}^{T}$, $C^{\prime}_{2z}$ symmetry implies $\psi_{-\mathbf{k}}=\begin{bmatrix}a_{\mathbf{k}}&c_{\mathbf{k}}&b_{\mathbf{k}}\end{bmatrix}^{T}$. Here, $a,b$ and $c$ are coefficients in our three-band basis, obtained by diagonalizing the mean-field Hamiltonian. They can be chosen to be real since $\chi_{\pm}$ are real. This implies that for the symmetric ($\delta\chi_{+}=\delta\chi_{-}$) superfluid Goldstone mode $\alpha$ is an even function: $\alpha_{s}(\mathbf{k},\mathbf{k}^{\prime})=\alpha_{s}(-\mathbf{k},-\mathbf{k}^{\prime})$, and for the anti-symmetric ($\delta\chi_{+}=-\delta\chi_{-}$) phonon mode, $\alpha$ is an an odd function: $\alpha_{p}(\mathbf{k},\mathbf{k}^{\prime})=-\alpha_{p}(-\mathbf{k},-\mathbf{k}^{\prime})$, as quoted in the main text. Finally, we integrate out the Goldstone modes to obtain an effective interaction term between the low-energy fermions:
$\frac{1}{2A\rho_{G}}\sum_{\mathbf{k},\mathbf{k}^{\prime},\mathbf{q}}f^{\dagger}_{\mathbf{k}+\mathbf{q}}f_{\mathbf{k}}f^{\dagger}_{\mathbf{k}^{\prime}-\mathbf{q}}f_{\mathbf{k}^{\prime}}\frac{\alpha_{G}(\mathbf{k}+\mathbf{q},\mathbf{k})\alpha_{G}(\mathbf{k}^{\prime}-\mathbf{q},\mathbf{k}^{\prime})}{q_{x}^{2}+\kappa q_{y}^{2}}.$ (A.22)
@@
We note that the strength of the interaction approaches a finite limit as $\mathbf{q}\rightarrow 0$, analogous to the case of typical Goldstone modes such as acoustic phonons and gapless spin waves [73], since $\alpha(\mathbf{k}+\mathbf{q},\mathbf{k})$ is linear in $\mathbf{q}$ for small $\mathbf{q}$. Let us define the Goldstone mediated interactions strength:
$V_{G}(\mathbf{k},\mathbf{k}^{\prime},\mathbf{q})=\frac{1}{\rho_{G}}\frac{\alpha_{G}(\mathbf{k}+\mathbf{q},\mathbf{k})\alpha_{G}(\mathbf{k}^{\prime}-\mathbf{q},\mathbf{k}^{\prime})}{q_{x}^{2}+\kappa_{G}q_{y}^{2}}.$ (A.23)
@@
$\frac{1}{2A}\sum_{\mathbf{k},\mathbf{l}}f^{\dagger}_{\mathbf{l}}f^{\dagger}_{-\mathbf{l}}f_{-\mathbf{k}}f_{\mathbf{k}}V_{G}(\mathbf{k},-\mathbf{k},\mathbf{l}-\mathbf{k}).$ (A.24)
Let us define $V_{BCS}(\mathbf{k},\mathbf{l})=V_{G}(\mathbf{k},-\mathbf{k},\mathbf{l}-\mathbf{k})/A$. We plot $V_{BCS}$ this interaction for $l_{x}\rightarrow k_{x},l_{y}=k_{y}$ in Fig. A.4, to demonstrate the attractive/repulsive nature of the interactions mediated by $\delta\chi_{p/s}$. This motivates us to ask whether the Goldstone mode mediated interaction can lead to superconductivity. Decoupling the interaction term in the BCS channel, we get the following BdG Hamiltonian.
$H_{\text{BdG}}=\sum_{\mathbf{k}}\begin{pmatrix}f^{\dagger}_{\mathbf{k}}&f_{-\mathbf{k}}\end{pmatrix}\begin{pmatrix}\epsilon_{\mathbf{k}}&\Delta_{BCS}(\mathbf{k})\\ \Delta^{*}_{BCS}(\mathbf{k})&-\epsilon_{-\mathbf{k}}\end{pmatrix}\begin{pmatrix}f_{\mathbf{k}}\\ f^{\dagger}_{-\mathbf{k}}\end{pmatrix}$ (A.25)
@@
$\Delta_{BCS}(\mathbf{k})=\sum_{\mathbf{l}}V_{BCS}(\mathbf{k},\mathbf{l})\langle f_{-,-\mathbf{l}}f_{-,\mathbf{l}}\rangle$ (A.26)
$\epsilon_{\mathbf{k}}=\epsilon_{-\mathbf{k}}$ due to the $C^{\prime}_{2z}$ symmetry. For a range of parameters, we find self-consistent superconducting solutions, as plotted in Fig. 4 in the main text. The gap is maximized in the region where the interactions are strongly attractive.
While we focus on $\mathbf{Q}\neq 0$ states in this work, the above arguments let us draw conclusions about the (charge-imbalanced) $\mathbf{Q}=0$ exciton condensed state as well. In this case, translational symmetry is not broken and therefore, there is only one Goldstone mode from the breaking of $U(1)_{r}$ symmetry: the superfluid mode $\delta\chi_{s}$. The fermion quasiparticles can be represented in terms of the two bands $f_{c,\mathbf{k}}$ and $f_{v,\mathbf{k}}$. The coupling between the Goldstone mode and the fermion quasiparticles, as shown for the superfluid mode above, is an even function of momentum in this case as well due to the $C^{\prime}_{2z}$ symmetry. Following the discussion in the main text, this results in interactions that are repulsive and even-parity, ruling out superconductivity.
Figure A.4: Left/Right panel: Phonon/superfluid mode mediated interaction matrix elements for ${\mathbf{l}}\rightarrow{\mathbf{k}}$ in Eq. A.24. The parameters used are $n=-0.7\times 10^{12}$ cm-2, $V_{b}=12$ meV and a $59\times 59$ momentum mesh.
IV.2 Spiral X-DW
There is another possible exciton density wave phase where the system spontaneously breaks $C^{\prime}_{2z}$ symmetry and picks either $\mathbf{Q}$ or $\mathbf{Q}$ for the exciton ordering. In this “spiral exciton density wave phase”, the phase of the exciton order parameter forms a spiral pattern. However, in our system, this is not the energetically preferred mean-field state. In this phase, a combination of the generators remains as a symmetry,for example, the $U(1)_{r}$ phase transformation, $\chi=\chi_{+}e^{i(\mathbf{Q}\cdot\mathbf{r})}\rightarrow\chi_{+}e^{i(\mathbf{Q}\cdot\mathbf{r}+\theta)}$, can be compensated by a lattice translation with $\alpha=-\theta$. This compensation entails the breaking of effectively one symmetry generator leading to one Goldstone mode.
For completness, we also examine the situation of Goldstone mediated superconductivity in the spiral X-DW. The Hubbard-Stratonovich field associated with the sprial X-DW is $\chi_{\mathbf{r}}=e^{i\textbf{Q}\cdot\mathbf{r}}(\chi+\delta\chi_{\mathbf{r}})$ where the spiral X-DW has a periodicity given by Q, $\chi$ is the uniform mean field exciton condensate, and $\delta\chi_{\mathbf{r}}$ is the fluctuation. We note that we have extracted out the fast oscillatory behavior (arising from Q) explicitly in the order parameter. The corresponding momentum space mean-field order parameter is self-consistently determined from the gap equation (which follows from Eq. A.11),
$\chi=\frac{U_{0}}{A}\sum_{\mathbf{k}^{\prime}}\left(n_{F}(\epsilon_{-,\mathbf{k}})-n_{F}(\epsilon_{+,\mathbf{k}})\right)\frac{\chi}{2E_{\mathbf{k}^{\prime}}}.$ (A.27)